OpenStudy (soupp):

I have x+y=49 .10x+.25y=7.45

jigglypuff314 (jigglypuff314):

you can use the substitution method to solve this :) x+y=49 -> y = -x + 49 plug into 0.1x + 0.25y = 7.45 to get 0.10x + 0.25(-x+49) = 7.45 then distribute, simplify, and solve for x :)